# Core Context Aware Transformers for Long Context Language Modeling

### v2

@@ 3.2 Globality-aware Pooling Module
7:  $\mathbf{K}^{\text{G}}{=}{\bf C}{\bf W}^{K}$, $\mathbf{V}^{\text{G}}{=}{\bf C}{\bf W}^{V}$ // Globality-aware Pooling Module
8:  $\mathbf{K}^{\text{L}}=\mathbf{X}\mathbf{W}^{K}$, $\mathbf{V}^{\text{L}}=\mathbf{X}\mathbf{W}^{V}$ // Locality-preserving Module
@@ Acknowledgments
This work was partially supported by the Joint Funds of the National Natural Science Foundation of China (Grant No.U24A20327), Key-Area Research and Development Program Guangdong Province 2018B010107001, the Major Key Project of Peng Cheng Laboratory (PCL) PCL2023A08, and TCL Science and Technology Innovation Fund, China.
@@
Our work addresses the critical challenge of computational inefficiency in long-context modeling for large language models. By introducing a core context aware attention mechanism, we achieve a reduction in attention complexity to linear time while preserving essential semantic interactions. This breakthrough enables LLMs to process contexts up to 128K tokens with a 7.9$\times$ speedup over full self-attention, without compromising accuracy. This promotes a more environmentally friendly and sustainable approach to AI development, aligning with the growing demand for greener and more energy-efficient technologies in the field of AI.
@@ C.5 More Ablation Studies
Effect of Group Size $g$. To investigate the effect of different group sizes $g$, we implement the proposed CCA-Attention with different $g$in $\{2,4,8,16,32,64\}$. In Table  8, as $g$ increases, the computational efficiency improves while the PPL increases. Upon closer examination, the smallest group size $g$ captures the most comprehensive information, which translates to the highest computational cost but also the optimal PPL. Conversely, an excessively large $g$ leads to an overemphasis on globality-aware attention, compressing information to the point where crucial semantic nuances may be overlooked, thereby curtailing performance. To strike a balance between computational efficiency and model performance, we have selected $g\small{=}16$ as the default training setting.
@@
Effect of Local Window Size $s$. To systematically evaluate the influence of different local window sizes $w$, we implement the proposed CCA-Attention across a range of $s$in $\{256,512,1024,2048,4096\}$. In Table 9, an increase in $s$ correlates with lower PPL, but this is counterbalanced by a rise in computational cost. A larger $s$ captures more contextual information with neighborhood tokens, but also increases computational demands. Conversely, a smaller $s$, indicative of a limited receptive field, constrains the exchange of information within the locality-preserving module, resulting in diminished performance. Striking a balance between computational efficiency and model efficacy, we opted for $s\small{=}1024$ as the default training setting.

### v3

@@ 3.2 Globality-aware Pooling Module
[…]
@@ Acknowledgments
This work was partially supported by the Joint Funds of the National Natural Science Foundation of China (Grant No.U24A20327), Key-Area Research and Development Program Guangdong Province 2018B010107001, the Major Key Project of Peng Cheng Laboratory (PCL) PCL2023A08, Postdoctoral Fellowship Program of CPSF (GZC20251043), and TCL Science and Technology Innovation Fund, China.
@@
Our work addresses the critical challenge of computational inefficiency in long-context modeling for large language models. By introducing a core context aware attention mechanism, we achieve a reduction in attention complexity to linear time while preserving essential semantic interactions. This enables LLMs to process contexts up to 128K tokens with a 7.9$\times$ speedup over full self-attention, without compromising accuracy. This promotes a more environmentally friendly and sustainable approach to AI development, aligning with the growing demand for more energy-efficient AI.
@@ C.5 More Ablation Studies
[…]
@@
[…]